Calculate the pH of a solution that is 0.0060 M in HClO4 and 0.0060 M in HCN. Ka of HCN is 6.2 × 10−10.
Answer #1

Since HClO4 is a strong acid, it is completely dissociated to form H+(aq) and ClO4-(aq).

Hence [H+(aq)] due to HClO4 = [HClO4] = 0.0060 M

Since HCN is a very weak acid with a very small value of Ka, H+(aq) formed due to its dissociation is very small.

Hence we can neglect [H+(aq)] due to dissociation of HCN.

Hence total [H+(aq)] = 0.0060 M

=> pH = - log[H+(aq)] = - log[0.0060] = 2.22 (Answer)